OpenStudy (anonymous):

Sandy is Working with a carpenter to frame a house. They are using 8 foot long boards, but each board must be cut to be 7 feet, 10 3/4 inches long. How much is cut off each board? A. 3/4 B. 1 1/4 C. 1 3/4 D. 7

OpenStudy (anonymous):

\[\large 8 \times 12 ~=~ 96\\\large 7 \times 12 ~=~ 84~+~10\frac{3}{4}~=~94\frac{3}{4}\\\large 96~-~94\frac{3}{4}~=~?\]